Hi I was wondering if somebody could help me find the equation of a line perpendicular to y=-5x+1 and passing through (-10,3).
Answer by stanbon(75887)

perpendicular to y=-5x+1 and passing through (-10,3).
---
All lines perpendicular to that line have slope m = 1/5
----
Find the intercept using: y = mx + b
3 = (1/5)*-10 + b
3 = -2 + b
b = 5
---------------
Equation you want:
y = (1/5)x + 5
====================
cheers,
Stan H.